OpenStudy (anonymous):

Evaluate. P(7,2) A.5 B.120 C.21 D.42

OpenStudy (whpalmer4):

\[P(n,r) = \frac{n!}{(n-r)!}\]Do you know how to evaluate factorials?

OpenStudy (anonymous):

I forgot. :\

OpenStudy (whpalmer4):

Well, \[n! = n*(n-1)*(n-2)*...(2)(1)\] so \[5! = 5*4*3*2*1\]etc.

OpenStudy (anonymous):

okay

OpenStudy (whpalmer4):

I would advise you take a moment to think before multiplying and dividing all those numbers :-) For example: \[\frac{5!}{3!} = \frac{5*4*3*2*1}{3*2*1} = \frac{5*4\cancel{*3*2*1}}{\cancel{3*2*1}} = 5*4\] I think you'll agree that that is easier to compute than multiplying and dividing the whole bunch, right?

OpenStudy (anonymous):

Okay let me see if i can figure out the answer

OpenStudy (anonymous):

Is it 42?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (aravindg):

Yes it is right :)

OpenStudy (anonymous):

Thanks

OpenStudy (whpalmer4):

Yep, \[\frac{7!}{(7-2)!} = \frac{7!}{5!} = \frac{7*6*5!}{5!} = \frac{7*6\cancel{*5!}}{\cancel{5!}} = 7*6 = 42\]